OpenStudy (anonymous):

For what value of x is sin x = cos 19°, where 0°< x < 90°? 38° 71° 26° 19°

OpenStudy (anonymous):

@TheSmartOne @5sauselover21

OpenStudy (anonymous):

@LeilaJudeh

OpenStudy (freckles):

think co function identity

OpenStudy (freckles):

\[\sin(90^o-x)=\cos(x)\]

OpenStudy (anonymous):

still confused......

OpenStudy (freckles):

compare cos(x) to cos(19) what must x be?

OpenStudy (campbell_st):

look at it this way  the sin of what angle given a/h

OpenStudy (anonymous):

im thinking A or C?

OpenStudy (anonymous):

nevermind figured it out! its 71 :)

TheSmartOne (thesmartone):

Correct, if you did it @freckles 's way then: \(\color{blue}{\text{Originally Posted by}}\) @freckles \[\sin(90^o-x)=\cos(x)\] \(\color{blue}{\text{End of Quote}}\) \(\sf sin(90^o-x)=cos(19)\) So x=19 sin (90-19) = cos(19) sin (71) = cos(19) So the angle for sin that equals cos 19 is sin 71 Good job :P

OpenStudy (freckles):

 was the point of @campbell_st 's post his post basically was explaining the identity

OpenStudy (freckles):

I mean which is probably way better than me just stating the identity :p